### v1

# Continual Segmentation with Disentangled Objectness Learning and Class Recognition

Abstract
Most continual segmentation methods tackle the problem as a per-pixel classification task. However, such a paradigm is very challenging, and we find query-based segmenters with built-in objectness have inherent advantages compared with per-pixel ones, as objectness has strong transfer ability and forgetting resistance. Based on these findings, we propose CoMasTRe by disentangling continual segmentation into two stages: forgetting-resistant continual objectness learning and well-researched continual classification. CoMasTRe uses a two-stage segmenter learning class-agnostic mask proposals at the first stage and leaving recognition to the second stage. During continual learning, a simple but effective distillation is adopted to strengthen objectness. To further mitigate the forgetting of old classes, we design a multi-label class distillation strategy suited for segmentation. We assess the effectiveness of CoMasTRe on PASCAL VOC and ADE20K. Extensive experiments show that our method outperforms per-pixel and query-based methods on both datasets. Code will be available at https://github.com/jordangong/CoMasTRe.
1 Introduction
Neural nets have dominated computer vision, from coarse-grained tasks like image classification to fine-grained tasks like segmentation and detection. However, the success is only sound for single-shot learning, i.e., training once to perform well with fully annotated datasets without considering the later learning process. Continual learning aims to mimic human learning by gradually obtaining knowledge in a sequential fashion [15, 17]. Particularly, continual learning shines in dense prediction tasks, such as semantic segmentation [3, 12, 5], since the annotations are laborious to obtain, and sometimes the learned samples are inaccessible, such as the nature of stream data in autonomous driving [32] or patient privacy in medical imaging [33].
(a) Mask proposal transferred to unseen classes.
(b) Mask proposals after finetuning on new classes.
Figure 1: Hidden properties inside query-based segmenters. Objectness in query-based methods helps generalize mask proposals on unseen classes similar to learned classes (top). Additionally, because of the transfer ability of objectness, query-based methods are resistant to catastrophic forgetting of mask proposals (bottom).
However, machines tend to forget the early tasks if we train them to learn the new ones, also termed as catastrophic forgetting [16], which is the critical obstacle toward continual learning. Moreover, dense prediction tasks tackled as per-pixel classification or regression by analyzing patterns around each pixel are inherently difficult, which makes the continual learning of these tasks even more challenging. Could we mitigate forgetting in continual segmentation in a more feasible way? The answer lies in a new paradigm: mask classification, streamlining continual segmentation as learning without forgetting of class-agnostic binary masks and class recognition, via query-based segmenters [9, 45, 10, 52]. Different from conventional per-pixel-based segmenters [8, 40, 50], from our observation, query-based segmenters have strong objectness (the ability to find objects in images) built-in, which is beneficial to continual segmentation for two reasons. Firstly, as the background is not silenced when learning to propose masks, the objectness can be transferred to unseen classes. As illustrated in Fig. 0(a), compared to DeepLabv3 [8], a per-pixel segmentation baseline, query-based Mask2Former [10] proposes substantially better masks on unseen classes if the model has learned similar classes before, e.g., learning vehicles, such as planes, boats, buses, and cars, facilitates the later learning of trains. Secondly, we observe that the objectness alleviates the forgetting of old class mask proposals (see Fig. 0(b)), as the model can still propose comparable masks on old classes, even finetuning on new classes.
Based on these findings, we propose Continual Learning with Mask-Then-Recognize TRansformer decoder, termed as CoMasTRe, for continual segmentation. CoMasTRe disentangles the segmentation task to objectness learning and class recognition by learning to propose class-agnostic masks at the first stage and leaving recognition to the second stage. In continual settings, since CoMasTRe inherits the objectness of query-based methods, the learning of new mask proposals is considerably eased. This design also simplifies continual segmentation to forgetting-resistant continual objectness learning and a continual classification task. Specifically, CoMasTRe mitigates forgetting in two folds: (i) a simple but effective objectness distillation to reinforce old class objectness during long learning processes, and (ii) a multi-label class distillation strategy with task-specific classifiers to alleviate old class forgetting. For evaluation, we conduct extensive experiments on PASCAL VOC 2012 and ADE20K, showcasing that CoMasTRe reaches a new state-of-the-art on both datasets. In particular, we boost incremented class performance compared with previous best methods on VOC. Additionally, CoMasTRe significantly outperforms prior arts in all classes on ADE20K.
To sum up, the contributions of this paper are as follows:
  • To leverage the properties of objectness, we propose Continual learning with Mask-Then-Recognize TRansformer decoder, termed as CoMasTRe, to simply continual segmentation by decoupling the task into objectness learning and class recognition.
  • To tackle the forgetting issue, we strengthen the objectness using a simple but effective distillation strategy in the first stage and preserve class knowledge in the second stage with task-specific classifiers and multi-label class distillation tailored to segmentation.
  • Through extensive benchmarking on two datasets, our method outperforms prior query-based methods by up to $2.1\%$ on ADE20K and surpasses per-pixel methods on new classes by up to $32.16\%$ on PASCAL VOC.
2 Related Work
Query-based Image Segmentation. Query-based segmenters [9, 10, 52, 21] have unified image segmentation via mask classification and solve semantic, instance, and panoptic segmentation using the same framework. However, the mask classification paradigm was originated from query-based detectors, such as DETR [2] and its variants [61, 24, 24, 54], where we train the model to predict proposals and class labels with learnable queries. MaskFormer [9] was the first to introduce mask classification in semantic segmentation, which was dominated by per-pixel classification paradigm via convolutional nets [8, 49, 44, 57] or Transformers [57, 40, 50]. Following MaskFormer, Mask2Former [10] introduced multiple improvements, such as masked attention for mask refinement and using multi-scale features for better small object performance. To speed up the convergence, $k$MaX-DeepLab [52] generated queries from clustered image features. Recently, OneFormer [21] provided a generalist solution by joint training of all segmentation tasks. However, all of these segmenters learn to predict to mask and class simultaneously and are not suitable for decoupled continual segmentation.
Continual Segmentation. Learning new knowledge by fintuning the model is a native way, but resulting catastrophic forgetting of old knowledge [16, 29, 37]. Continual learning aims to alleviate the forgetting issue by balancing the learning of new knowledge (plasticity) and the preserving of the old one (stability). While the research focus of continual learning is on classification [22, 25, 36, 1, 11, 6, 47, 39], continual segmentation started to thrive recently [3, 12, 30, 7, 34, 53, 55, 5, 38, 48, 60, 56]. As pointed out in MiB [3], the background shift in continual segmentation aggravates the forgetting issue. Cermelli et al. [3] proposed unbiased distillation and classification to mitigate the background shift. PLOP [12] circumvented the background shift issue by feature distillation and pseudo-labeling. Based on PLOP, REMINDER [34] introduced class-weighted knowledge distillation to discriminate the new and old classes. With the same motivation, Incrementer [38] designed a class deconfusion strategy and achieved competitive performance via a Transformer-based segmenter. Recently, RL-Replay [60] set a strong replay-based baseline via reinforcement learning. Some works [4, 20] started to focus on more realistic scenarios by integrating weak supervision with continual learning. The methods above follow the per-pixel classification paradigm. Although CoMFormer [5] was the first to present mask classification in continual segmentation, it applied a standard framework with distillation and pseudo-labeling and did not leverage the benefit of objectness. In contrast, our CoMasTRe exploits objectness and decouples the continual segmentation into forgetting-resistant objectness learning and class recognition.
Continual Dynamic Networks. In addition to distillation methods appearing in continual segmentation, dynamic networks with parameter expansion [51, 43, 59, 13, 47, 46, 39, 39, 35] also play a key role in continual learning. DER [51] is an early work on the dynamic structure by duplicating the backbone for each task to mitigate the forgetting issue. However, DER suffers from the explosion of parameters after a long learning process. FOSTER [43] relieved the problem through a model compression stage. MEMO [59] went further by decoupling the network and only expanding specialized layers. Recently, more works [13, 47, 46, 39, 39, 35] started to leverage the dynamic nature of Transformers. DyTox [13] proposed a dynamic token expansion mechanism by learning a task-specialized prompt at each step. The concurrent work, L2P [47], started a new fashion of continual learning: learning to prompt pretrained models and extracting task-specialized features. Thanks to our Transformer-based decoder, we adopt task queries and task-specific classifiers in our class decoder to further alleviate forgetting.
3 Method
3.1 Problem Definition
Image segmentation has been unified as a mask classification problem, i.e., to propose class-agnostic masks and predict the corresponding class labels simultaneously. Formally, we define the dataset $\mathcal{D}$ containing pairs of image and target, and each sample pair contains an image $x\in\mathbb{R}^{C\times H\times W}$ and its target $y$, where $C$, $H$ and $W$ denote channel number, height and width of the image, respectively. Each target $y$ is composed by $M$ ground truth (GT) binary masks and class labels, denoted $y=\{(m_{i}^{\mathrm{gt}},c_{i}^{\mathrm{gt}})\mid m_{i}^{\mathrm{gt}}\in\{0,1\}^{H\times W},c_{i}^{\mathrm{gt}}\in\mathcal{C}\}_{i=1}^{M}$, where $\mathcal{C}$ is the class label space. Unlike per-pixel classification methods, background class is excluded from annotation.
In terms of continual segmentation, the model sequentially learns to predict the masks of new classes but not to forget old ones. Formally, we divide the learning process to $T$ tasks, and at each step $t=1,2,\dots,T$ we train the model to predict a unique set of classes $\mathcal{C}^{t}$, where $\bigcap_{t=i}^{T}\mathcal{C}^{t}=\varnothing$ and $\bigcup_{t=i}^{T}\mathcal{C}^{t}=\mathcal{C}$. Only the part of the annotations containing new classes $\mathcal{C}^{t}$ are provided at step $t$, while the model should be able to predict all learned classes $\mathcal{C}^{1:t}$.
3.2 CoMasTRe Architecture
Figure 2: CoMasTRe Architecture. $\bigotimes$ denotes the dot product between positional embeddings $\mathcal{E}_{\mathrm{pos}}$ and pixel embeddings $\mathcal{E}^{4}_{\mathrm{pixel}}$. CoMasTRe uses a two-stage image segmenter including three components: (a) a backbone and a pixel decoder producing pixel embedding, (b) a mask decoder $f$ with learnable positional queries $\mathcal{Q}_{\mathrm{pos}}$ for objectness learning, and (c) a class decoder $g$ with a set of task queries $\mathcal{Q}_{\mathrm{task}}$ for class recognition.
Thanks to the properties of objectness (see Fig. 1) and widely studied continual classification, we argue that solving the continual segmentation problem with mask classification is inherently advantageous. Therefore, we propose CoMasTRe, which adopts Mask2Former [10] meta-architecture, keeping the backbone and the pixel decoder while adding two newly designed Transformer decoders to help disentangle objectness learning (stage 1) and class recognition (stage 2), as shown in Fig. 2. The overall training process of our CoMasTRe can be summarized as:
  1. 1.
    When an image comes to CoMasTRe, the backbone, and the pixel decoder first encode and then decode the image to pixel embeddings. We take out the pixel embeddings from all the layers of the pixel decoder for the following process as $\{\mathcal{E}_{\mathrm{pixel}}^{l}\}^{4}_{l=1}$, where $l$ denotes different layers and there are 4 layers in total (Fig. 2 (a)).
  2. 2.
    Next, we randomly initialize learnable positional queries $\mathcal{E}_{\mathrm{pos}}$ and input them to the objectness learning stage (stage 1), to extract positional embeddings $\mathcal{E}_{\mathrm{pos}}$ through the mask decoder $f$ for predicting class-agnostic mask proposals $\mathcal{M}$ and objectness scores $\mathcal{S}$. The training involves bipartite matching with ground truth (Fig. 2 (b)).
  3. 3.
    After that, matched positional embeddings $\mathcal{E}_{\mathrm{pos}}^{\mathrm{m}}$ from stage 1 with pixel embeddings are input to class decoder $g$ for recognition (stage 2). To reduce task interference during continual learning, task queries $\mathcal{Q}_{\mathrm{task}}$ and task-specific classifiers $\phi_{1},\dots,\phi_{t}$ are learned to specialize the class knowledge of each task. (Fig. 2 (c)).
  4. 4.
    Finally, we obtain segmentation results by combining the mask proposals and objectness scores from stage 1 with the class prediction from stage 2.
The details of the two stages will be illustrated in Sec. 3.2.1 and Sec. 3.2.2.
3.2.1 Stage 1: Objectness Learning
As illustrated in Fig. 2 (b), our objectness learning mainly relies on a mask decoder. The mask decoder $f$ is composed of $L$ blocks of Transformer layers [42], taking $N$ learnable positional queries $\mathcal{Q}_{\mathrm{pos}}=\{q_{1},\dots,q_{N}\}\in\mathbb{R}^{N\times d}$ and intermediate pixel embeddings $\{\mathcal{E}^{l}_{\mathrm{pixel}}\}_{l=1}^{3}$ as input. Then, it outputs the positional embeddings $\mathcal{E}_{\mathrm{pos}}=\{e_{1},\dots,e_{N}\}\in\mathbb{R}^{N\times d}$ for mask proposals, objectness scores, and stage 2 recognition. The mask proposals =$\mathcal{M}=\{m_{1},\dots,m_{N}\}\in[0,1]^{N\times H\times W}$ is computed as $\mathcal{M}=\mathrm{sigmoid}(\mathrm{Upsample}(\mathrm{MLP}(\mathcal{E}_{\mathrm{pos}})\cdot\mathcal{E}_{\mathrm{pixel}}^{4}))$, where the $\mathrm{MLP}(\cdot)$ acts as a non-linear transformation function and $\mathrm{Upsample}(\cdot)$ interpolates the logits to image size. Note here each positional embedding produces a mask proposal. At the same time, $\mathcal{E}_{\mathrm{pos}}$ is input to our objectness head, in which a binary classifier is designed to output the objectness score $\mathcal{S}=\{s_{1},\dots,s_{N}\}\in[0,1]^{N}$ to indicate whether the mask proposals containing objects or not. During training, we first perform bipartite matching considering the cost between $N$ mask proposals $\{m_{i}\}_{i=1}^{N}$ and $M$ ground truth masks $\{m_{j}^{\mathrm{gt}}\}_{j=1}^{M}$. Here, we assume $N\gg M$ and pad the ground truth with “no object” $\varnothing$ to ensure a one-to-one matching between predictions and ground truth. After solving the matching via the Hungarian algorithm [23], we get the optimal permutation of $N$ elements, denoted as $\sigma$. Then, we optimize this stage using ground truth masks:
$\mathcal{L}_{\mathrm{obj}}=\sum_{j=1}^{N}\big{[}$ $-\mathds{1}_{m_{j}^{\mathrm{gt}}\neq\varnothing}{\log{s_{\sigma_{j}}}}-\mathds{1}_{m_{j}^{\mathrm{gt}}=\varnothing}{\log{(1-s_{\sigma_{j}})}}$ (1)
$+\mathds{1}_{m_{j}^{\mathrm{gt}}\neq\varnothing}\mathcal{L}_{\mathrm{mask}}(m_{\sigma_{j}},m_{j}^{\mathrm{gt}})\big{]},$
where $\mathcal{L}_{\mathrm{mask}}$ is the sum of binary cross-entropy loss and Dice loss [41]. Based on the matching results, $M$ matched positional embeddings $\mathcal{E}^{\mathrm{m}}_{\mathrm{pos}}=\{e_{j}\mid e_{j}\in\mathcal{E}_{\mathrm{pos}},m_{j}^{\mathrm{gt}}\neq\varnothing\}$ are traced back and fed forward to stage 2 class decoder. Here, we supervise the objectness scores by encouraging the matched ones and suppressing the unmatched ones, otherwise, the model tends to produce high objectness scores for all the proposals. The remaining unmatched embeddings $\mathcal{E}^{\mathrm{u}}_{\mathrm{pos}}=\mathcal{E}_{\mathrm{pos}}\setminus\mathcal{E}^{\mathrm{m}}_{\mathrm{pos}}$ are reserved for distillation to alleviate forgetting (see Sec. 3.3).
3.2.2 Stage 2: Class Recognition
Our stage 2 aims to recognize the object inside the mask proposals. Thus, we use matched positional embeddings along with pixel embeddings as input for class recognition. Besides, inspired by DyTox [13], we integrate task queries with matched positional embeddings as input to task-specific classifiers in the class decoding process, which facilitates continual learning by reducing task interference, resulting in more specialized classifiers.
Since continual segmentation contains $T$ tasks, to illustrate our class recognition stage, we assume the current learning step is $t$. We denote task queries $\mathcal{Q}_{\mathrm{task}}=\{\theta_{1},\dots,\theta_{t}\}$. As shown in Fig. 2 (c), for each positional embedding $e\in\mathcal{E}_{\mathrm{pos}}^{\mathrm{m}}$ and each task query $\theta\in\mathcal{Q}_{\mathrm{task}}$, we take $(e,\theta)$ as a pair and feed it with pixel embedding to the class decoder $g$ (a Transformer block shared across tasks), obtaining the localized task embedding $k$. For each positional embedding, we obtain its task embeddings $\mathcal{E}_{\mathrm{task}}=\{k_{1},\dots,k_{t}\}$. Then, $t$ task-specific classifiers $\Phi^{t}=\{\phi_{1},\dots,\phi_{t}\}$ are applied to $\mathcal{E}_{\mathrm{task}}$, where each classifier $\phi_{j}\in\Phi^{t}$ is specialized to predict the classes in this task $\mathcal{C}^{j}$, i.e., $\phi_{j}:\mathbb{R}^{d}\mapsto{\mathcal{C}^{j}}$. In this way, we obtain the class probability of each proposal as $p=\mathrm{sigmoid}([\phi_{1}(k_{1}),\dots,\phi_{t}(k_{t})])$, as each positional embedding corresponds to a mask proposal. By feeding all the positional embeddings, we get the class probability of all proposals $\mathcal{P}=\{p_{i}\}_{i=1}^{M}$. We supervise this stage with classification loss using ground truth class labels:
$\mathcal{L}_{\mathrm{cls}}=-\frac{1}{M}\sum_{i=1}^{M}\sum_{c\in\mathcal{C}^{t}}\big{[}$ $\mathds{1}_{c=c_{i}^{\mathrm{gt}}}(1-p_{i,c})^{\gamma}\log{p_{i,c}}$ (2)
$+$ $\mathds{1}_{c\neq c_{i}^{\mathrm{gt}}}p_{i,c}^{\gamma}\log{(1-p_{i,c})}\big{]},$
where $p_{i,c}$ is the predicted probability of class $c$ for the $i$-th proposal, $\gamma$ is a hyperparameter. Here, we use focal loss [27] instead of cross-entropy loss for better calibration [31], which smooths the distillation process during continual learning. The overall segmentation loss is the sum of objectness loss and classification loss, i.e., $\mathcal{L}_{\mathrm{seg}}=\mathcal{L}_{\mathrm{obj}}+\mathcal{L}_{\mathrm{cls}}$. We jointly train two stages by minimizing $\mathcal{L}_{\mathrm{seg}}$. During inference, we set an objectness threshold $\alpha$ at stage 1 to filter low confident prediction and feed high objectness embedding to stage 2 for mask recognition.
3.3 Learning without Forgetting with CoMasTRe
Learning new classes by naively finetuning the model will cause catastrophic forgetting. To alleviate this issue, our CoMasTRe separately considers the distillation of both stages. In this way, predictions on base classes can also be preserved when learning new classes.
3.3.1 Objectness Distillation
As shown in Fig. 3 (a), we first train the model with bipartite matching and get matched positional embeddings and unmatched ones. However, the objectness scores of unmatched ones quickly diminish as the ground truth does not contain masks from old classes, resulting in incorrect objectness scores on old classes. Thus, we introduce objectness score distillation with $\mathcal{L}_{\mathrm{os-kd}}$ to preserve objectness scores learned in previous steps. Moreover, in spite of the robust nature of objectness, could we further improve the objectness during continual learning? To answer this question, we propose another two distillation losses $\mathcal{L}_{\mathrm{mask-kd}}$ and $\mathcal{L}_{\mathrm{pe-kd}}$ for mask proposals and position embeddings, respectively (see Fig. 3 (b)). Following previous methods [25, 3], we distill the outputs from the last step model to mitigate the objectness forgetting on old classes.
Formally, we denote unmatched mask proposals as $\mathcal{M}^{\mathrm{u}}=\{m^{\mathrm{u}}_{1},\dots,m^{\mathrm{u}}_{N-M}\}$ and objectness scores as $\mathcal{S}^{\mathrm{u}}=\{s^{\mathrm{u}}_{1},\dots,s^{\mathrm{u}}_{N-M}\}$, which are the corresponding outputs from unmatched positional embeddings $\mathcal{E}_{\mathrm{pos}}^{\mathrm{u}}=\{e^{\mathrm{u}}_{i},\dots,e^{\mathrm{u}}_{N-M}\}$. Additionally, the distillation process requires the corresponding outputs from the last step model, including mask proposals ~$\tilde{\mathcal{M}}^{\mathrm{u}}=\{\tilde{m}^{\mathrm{u}}_{1},\dots,\tilde{m}^{\mathrm{u}}_{N-M}\}$, objectness scores $\tilde{\mathcal{S}}^{\mathrm{u}}=\{\tilde{s}^{\mathrm{u}}_{1},\dots,\tilde{s}^{\mathrm{u}}_{N-M}\}$, and positional embeddings ~$\tilde{\mathcal{E}}_{\mathrm{pos}}^{\mathrm{u}}=\{\tilde{e}^{\mathrm{u}}_{1},\dots,\tilde{e}^{\mathrm{u}}_{N-M}\}$. We freeze the parameters of the last step model during distillation.
Distill objectness scores. We use vanilla knowledge distillation loss [19] to mitigate the diminishing of objectness scores by minimizing Kullback-Leibler (KL) divergence between the current unmatched objectness score $\mathcal{S}^{\mathrm{u}}$ and the last step ones $\tilde{\mathcal{S}}^{\mathrm{u}}$,
$\mathcal{L}_{\mathrm{os-kd}}=-\frac{1}{|\mathcal{S}^{\mathrm{u}}|}\sum_{i=1}^{|\mathcal{S}^{\mathrm{u}}|}\left[\tilde{s}_{i}^{\mathrm{u}}\log{\frac{s_{i}^{\mathrm{u}}}{\tilde{s}_{i}^{\mathrm{u}}}}+(1-\tilde{s}_{i}^{\mathrm{u}})\log{\frac{1-s_{i}^{\mathrm{u}}}{1-\tilde{s}_{i}^{\mathrm{u}}}}\right].$ (3)
As objectness scores play a key role during inference, this distillation is designed to maintain the objectness scores of previously learned mask proposals.
Distill mask proposals. To preserve the knowledge of the mask proposals but also focus on proposals with high objectness scores, we reweight mask distillation with objectness scores. By distilling masks with low objectness scores, we relax the case when objectness scores fail to indicate the objectness, as we observe some mask proposals generalize to unseen classes, but their objectness scores are low. The mask distillation loss is defined as
$\mathcal{L}_{\mathrm{mask-kd}}=\sum_{i=1}^{|\mathcal{M}^{\mathrm{u}}|}\frac{\omega_{i}}{\sum_{i=1}^{|\mathcal{M}^{\mathrm{u}}|}\omega_{i}}\mathcal{L}_{\mathrm{mask}}(m_{i}^{\mathrm{u}},\tilde{m}_{i}^{\mathrm{u}}),$ (4)
where $\omega_{i}=(\tilde{s}_{i}^{\mathrm{u}})^{\beta}$, and $\beta$ is a reweight strength hyperparameter.
Distill positional embeddings. Similar to mask distillation, we reweight position distillation when enforcing the unmatched positional embeddings $\mathcal{E}_{\mathrm{pos}}^{\mathrm{u}}$ to be similar to the last step ones ~$\tilde{\mathcal{E}}_{\mathrm{pos}}^{\mathrm{u}}$. Here, we encourage high cosine similarity between the current step ones and the last step ones, which assists in preserving positional information for later multi-label class distillation,
$\mathcal{L}_{\mathrm{pe-kd}}=\sum_{i=1}^{|\mathcal{E}_{\mathrm{pos}}^{\mathrm{u}}|}\frac{\omega_{i}}{\sum_{i=1}^{|\mathcal{E}_{\mathrm{pos}}^{\mathrm{u}}|}\omega_{i}}\left(1-\cos(e_{i}^{\mathrm{u}},\tilde{e}_{i}^{\mathrm{u}})\right),$ (5)
where $\cos(\cdot)$ means cosine similarity.
Finally, we distill the objectness by minimizing $\mathcal{L}_{\mathrm{obj-kd}}=\mathcal{L}_{\mathrm{os-kd}}+\mathcal{L}_{\mathrm{mask-kd}}+\mathcal{L}_{\mathrm{pe-kd}}$.
Figure 3: Learning without forgetting with CoMasTRe. To tackle catastrophic forgetting, CoMasTRe separated the distillation process into two stages, including objectness distillation and class distillation. We perform bipartite matching first, as in (a). Then, we distill the knowledge of objectness for remaining embeddings, as in (b). Finally, we select positional embeddings for the class distillation if they match with ground truth or have high objectness scores at the last step. The class knowledge is distilled from both matched and unmatched positional embeddings with a class decoder and task-specific classifiers, as in (c).
3.3.2 Class Distillation
As the segmentation is decoupled to a recognition task at the second stage, the continual learning of the second stage is essentially continual multi-label classification. We discuss class distillation in two scenarios: (i) distilling the class prediction from the last step high-objectness embeddings unmatched with GT, and (ii) distilling the class knowledge from matched embeddings (see in Fig. 3 (c)). During class distillation, we freeze the parameters of the last step model.
Distill from unmatched queries. This scenario is similar to the pseudo-labeling of the old classes but in a soft way, minimizing the KL divergence between the class probability of the current step and the last step. Given the last step unmatched embeddings ~$\tilde{\mathcal{E}}_{\mathrm{pos}}^{\mathrm{u}}$, we select $n$ embeddings with high objectness score ~$\tilde{\mathcal{E}}_{\mathrm{pos}}^{\mathrm{h}}=\{\tilde{e}_{i}^{\mathrm{u}}\mid\tilde{s}_{i}^{\mathrm{u}}>\alpha\}\in\mathbb{R}^{n\times d}$, where $\alpha$ is high objectness threshold. The corresponding current step embeddings $\mathcal{E}_{\mathrm{pos}}^{\mathrm{h}}$ are also selected. After class decoding, we obtain the last step class logits $\{\tilde{z}_{i}^{\mathrm{h}}\}_{i=1}^{n}$ and the current ones $\{z_{i}^{\mathrm{h}}\}_{i=1}^{n}$. The knowledge distillation loss is enforced on the probability of old classes $\mathcal{C}^{1:t-1}$, i.e., $\mathcal{L}_{\mathrm{cls-kd}}^{\mathrm{u}}=\frac{1}{n}\sum_{i=1}^{n}D_{\mathrm{KL}}(p_{i}^{\mathrm{h}}\|\tilde{p}_{i}^{\mathrm{h}})$, where $p_{i}^{\mathrm{h}}=\mathrm{softmax}(z^{\mathrm{h}}_{i})$ and $\tilde{p}^{\mathrm{h}}_{i}=\mathrm{softmax}(\tilde{z}^{\mathrm{h}}_{i})$ are the current step old class probability and the last step one decoded from $i$-th positional embedding. $D_{\mathrm{KL}}(\cdot)$ denotes KL divergence.
Distill from matched queries. For new samples, we consider its last step model prediction as the prior knowledge to mitigate forgetting. To extract the knowledge, we reuse the currently matched embeddings $\mathcal{E}_{\mathrm{pos}}^{\mathrm{m}}$ and get the old class probability activated by softmax $\{p_{i}^{\mathrm{m}}\}_{i=1}^{M}$. Additionally, we decode $\mathcal{E}_{\mathrm{pos}}^{\mathrm{m}}$ using the last step model, obtaining the last step probability $\{\tilde{p}_{i}^{\mathrm{m}}\}_{i=1}^{M}$ as the target distribution. During training, we minimize matched class distillation loss $\mathcal{L}_{\mathrm{cls-kd}}^{\mathrm{m}}=\frac{1}{M}\sum_{i=1}^{M}D_{\mathrm{KL}}(p_{i}^{\mathrm{m}}\|\tilde{p}_{i}^{\mathrm{m}})$.
In total, we preserve class knowledge by minimizing $\mathcal{L}_{\mathrm{cls-kd}}=\mathcal{L}_{\mathrm{cls-kd}}^{\mathrm{u}}+\mathcal{L}_{\mathrm{cls-kd}}^{\mathrm{m}}$. Besides, we follow prior classification works [51, 13] and use an auxiliary loss $\mathcal{L}_{\mathrm{cls-aux}}$ with a similar purpose to previous continual segmentation methods [34, 38]. It encourages more discriminative classification of similar classes by learning to predict a merged old class and new classes (1 + $|\mathcal{C}^{t}|$ classes in total) at each step. Together with segmentation loss, the loss of CoMasTRe is $\mathcal{L}=\mathcal{L}_{\mathrm{seg}}+\mathcal{L}_{\mathrm{kd}}+\mathcal{L}_{\mathrm{cls-aux}}$, where $\mathcal{L}_{\mathrm{kd}}=\mathcal{L}_{\mathrm{obj-kd}}+\mathcal{L}_{\mathrm{cls-kd}}$.
4 Experiments
4.1 Setup
Datasets. Following standard continual segmentation baselines [3, 12, 34], we evaluate semantic segmentation performance of CoMasTRe on PASCAL VOC 2012 [14] and ADE20K [58]. PASCAL VOC 2012 is a relatively small dataset with 20 object classes and background class, containing 10,582 samples for training and 1,449 for validation. ADE20K is a large-scale semantic segmentation dataset with 150 annotated classes, containing 20,210 and 2,000 samples for training and validation.
Table 1: Comparison with previous best methods on PASCAL VOC 2012 in mIoU (%). The highest and the second highest results are marked in bold and underline. $\ast$ means results from our re-implementation.
Paradigm Method 19-1 (2 tasks) 15-5 (2 tasks) 15-1 (6 tasks)
1-19 20 all avg 1-15 16-20 all avg 1-15 16-20 all avg
Per-Pixel MiB [3] 71.43 23.59 69.15 73.28 76.37 49.97 70.08 75.12 34.22 13.50 29.29 54.19
SDR [30] 68.52 23.29 66.37 71.48 75.21 46.72 68.64 74.32 43.08 19.31 37.42 54.52
PLOP [12] 75.35 37.35 73.54 75.47 75.73 51.71 70.09 75.19 65.12 21.11 54.64 67.21
REMINDER [34] 76.48 32.34 74.38 76.22 76.11 50.74 70.07 75.36 68.30 27.23 58.52 68.27
RCIL [53] 78.80 52.00 72.40 70.60 23.70 59.40
Joint 77.45 77.94 77.39 78.88 72.63 77.39 78.88 72.63 77.39
Query ${}^{\ast}$ [5] 75.35 24.06 72.91 75.46 74.68 48.47 68.44 72.97 48.97 23.28 48.18 64.16
Joint 77.09 71.39 76.82 78.60 71.12 76.82 78.60 71.12 76.82
CoMasTRe (ours) 75.13 69.51 74.86 76.66 79.73 51.93 73.11 75.97 69.77 43.62 63.54 70.63
Joint 78.57 68.32 78.08 80.83 69.28 78.08 80.83 69.28 78.08
Table 2: Comparison with previous methods on ADE20K in mIoU (%). The 1st and 2nd highest results are marked in bold and underline.
Paradigm Method 100-50 (2 tasks) 100-10 (6 tasks) 100-5 (11 tasks)
1-100 101-150 all avg 1-100 101-150 all avg 1-100 101-150 all avg
Per-Pixel MiB [3] 40.50 17.20 32.80 37.30 38.30 11.30 29.20 35.10 36.00 5.70 26.00 32.70
SDR [30] 40.52 17.17 32.79 37.31 37.26 12.13 28.94 34.48 33.02 10.63 25.61 33.07
PLOP [12] 41.87 14.89 32.94 37.39 40.48 13.61 31.59 36.64 35.72 12.18 27.93 35.10
REMINDER [34] 41.55 19.16 34.14 38.43 38.96 21.28 33.11 37.47 36.06 16.38 29.54 36.49
RCIL [53] 42.30 18.80 34.50 39.30 17.60 32.10 38.50 11.50 29.60
Joint 44.34 28.21 39.00 44.34 28.21 39.00 44.34 28.21 39.00
Query CoMFormer [5] 44.70 26.20 38.40 41.20 40.60 15.60 32.30 37.40 39.50 13.60 30.90 36.50
Joint 46.90 35.60 43.10 46.90 35.60 43.10 46.90 35.60 43.10
CoMasTRe (ours) 45.73 26.02 39.20 41.62 42.32 18.42 34.41 38.41 40.82 15.83 32.55 38.64
Joint 48.48 36.11 44.36 48.48 36.11 44.36 48.48 36.11 44.36
Continual Segmentation Protocols. Continual segmentation protocols include sequential, disjoint, and overlapped, where all of them split the segmentation into $T$ steps, and each step aims to learn a unique set of classes. The sequential is the easiest, as the training set contains the ground truth of each pixel. In contrast, the disjoint annotates past classes as background and excludes any training pixel belonging to future classes. The overlapped is the most challenging and realistic one, where the dataset only contains the ground truth of current classes, but both past and future classes are annotated as background. Following previous works [12, 34, 5], we adopt the overlapped in benchmarks.
We evaluate CoMasTRe in 6 settings, 3 on PASCAL VOC and 3 on ADE20K. Here, we use the $B$-$I$ notation in the paper, where $B$ and $I$ denote the number of base classes and incremented classes per step. Then, the number of total steps $T=1+(|\mathcal{C}|-B)/I$. For example, 15-1 on VOC means we start the learning process at 15 base classes, then continually train the model for 5 steps, one new class per step. In this way, the 3 settings on PASCAL VOC are (i) 19-1, 2-step learning starting with 19 classes and following 1 class, (ii) 15-5, 2-step learning starting with 15 classes and following 5 classes, and (iii) 15-1, consisting 6 learning steps with 15 classes at first and following 5 steps with one class each step. Similarly, the settings on ADE20K are 100-50, 100-10, and 100-5. Note that the longer the learning process, the harder it is to tackle forgetting.
Metric. We evaluate semantic segmentation performance using mean intersection over union (mIoU). In continual learning settings, the mIoU of base classes $\mathcal{C}^{1}$ and incremented classes $\mathcal{C}^{2:T}$ are also measured. The base class performance measures the stability of the model, while the incremented one reflects the plasticity. Additionally, following [12, 34, 5], we report the average mIoU of all learning steps as avg to evaluate the entire continual learning process. We denote joint as the joint training of all classes.
Implementation details. Following the previous continual segmentation benchmarks [3, 12, 5], we use ImageNet-pretrained ResNet101 [18] as the backbone. The pixel decoder in CoMasTRe is identical to the Mask2Former [10] one, the mask decoder consists of $L=9$ Transformer decoder blocks, and the class decoder consists of one block. We set the number of positional queries $N$ to 20 and 100 for PASCAL VOC and ADE20K, respectively. During training, we follow the Mask2Former [10] hyperparameters. We use AdamW [28] optimizer with an initial learning rate of $1\times 10^{-4}$ and weight decay of $0.05$ and apply a polynomial learning rate schedule. We train the model with a batch size of 16 for 20K iterations on PASCAL VOC and 160K iterations on ADE20K at the first learning step. In later steps, we half the learning rate to $5\times 10^{-5}$ and train for 1,000 iterations per class on VOC and 500 iterations per class on ADE20K. The images are augmented with random rescaling, flipping, and color jittering then cropped to $512\times 512$ resolution. We set reweight parameters $\gamma$ and $\beta$ to $2.0$. To evaluate the IoU of background during continual learning, we use the panoptic inference from Mask2Former [10]. During inference, we set the high objectness threshold $\alpha$ as $0.8$. Following CoMFormer [5], we use single-scale inference, and no replay is involved throughout training.
4.2 Quantitative Evaluation
We compare CoMasTRe with state-of-the-art continual segmentation methods on PASCAL VOC 2012 and ADE20K. Qualitative results are available in supplementary materials.
PASCAL VOC 2012. We compare our method with previous state-of-the-art methods on the 3 settings on PASCAL VOC: 19-1, 15-5, and 15-1 in Tab. 1. In general, CoMasTRe significantly outperforms query-based CoMFormer [5] and surpasses per-pixel baselines on incremented classes by a large margin. On 19-1, CoMasTRe shows off its strong plasticity in learning new classes, with 32.16 percent points (p.p) improvement while maintaining the knowledge of old classes. On 15-5, our model is on par with RCIL [53] on new classes while exceptionally great at maintaining old class performance, even surpassing the joint upper bound of per-pixel baselines by 0.85 p.p. 15-1 is much more difficult and involves 6 learning steps. However, CoMasTRe still beats RCIL with 4.14 p.p uplift on all classes. Overall, we strikes a balance between stability and plasticity.
ADE20K. As shown in Tab. 2, we report the results on ADE20K in 100-50, 100-10, and 100-5. In general, CoMasTRe outperforms previous state-of-the-art CoMFormer [5] in each setting by 0.8 p.p, 2.11 p.p, and 1.65 p.p on all classes, respectively. In 100-50, our method shows competitive performance with CoMFormer on incremented classes but with 0.8 p.p gain on all classes thanks to its higher upper bound. For longer learning processes, CoMasTRe reaches a comparable performance with REMINDER [34] on new classes but significantly boosts base class mIoU (3.36 p.p in 100-10 and 4.76 p.p in 100-5). Compared with CoMFormer, our CoMasTRe also learns better new classes with 2.82 p.p improvement in 100-10 and 2.23 p.p in 100-5 while preserving better old class knowledge, with 1.72 p.p and 1.32 p.p gains, respectively.
4.3 Ablation Studies
Joint training results. To resolve scalability concerns on mask-only matching, we demonstrate the joint performance of CoMasTRe, i.e., single-shot training on all classes. We compare the segmentation performance in mIoU on PASCAL VOC and ADE20k using ResNet50 (R50) and ResNet101 (R101) backbones in Tab. 3. The results show that CoMasTRe achieves comparable performance with Mask2Former and outperforms the previous continual model CoMFormer by $\sim$1.2 p.p on ADE20K.
Table 3: Semantic segmentation results mIoU in joint setting. Following continual settings, we use panoptic-style inference. The 1st and 2nd highest results are marked in bold and underline.
Method Backbone Datasets
VOC ADE20k
Mask2Former [10] R50 79.73 43.28
R101 80.04 43.51
CoMFormer [5] R50 76.01 42.58
R101 76.82 43.10
CoMasTRe (ours) R50 77.84 43.97
R101 78.08 44.36
Table 4: Objectness transfer ability results in PASCAL VOC 15-1 setting. For transfer learning, we continually train on VOC using a stage 1 mask decoder pretrained on COCO (w/o PASCAL classes). Otherwise, the model is trained on VOC only without pretraining.
Transfer 1-15 16-20 all avg
70.72 46.18 64.88 71.34
69.77 43.62 63.54 70.63
Table 5: Objectness distillation results in PASCAL VOC 15-1 setting, where $\mathcal{L}_{\mathrm{mask-kd}}$ for mask distillation, $\mathcal{L}_{\mathrm{os-kd}}$ for objectness score distillation, $\mathcal{L}_{\mathrm{pe-kd}}$ for position distillation.
$\mathcal{L}_{\mathrm{mask-kd}}$ $\mathcal{L}_{\mathrm{os-kd}}$ $\mathcal{L}_{\mathrm{pe-kd}}$ 1-15 16-20 all avg
13.14 35.42 18.44 47.43
63.75 37.97 57.61 62.72
66.23 38.57 59.64 64.61
69.77 43.62 63.54 70.63
Objectness transfer ability analysis. To study the transfer ability of objectness, we first pretrain a stage 1 mask decoder on modified COCO 2017 [26], where instances of the classes appearing in PASCAL VOC are removed to prevent information leaks. Then, we transfer the stage 1 and continually train it on 15 base classes of VOC with objectness distillation. Finally, we follow standard procedures by training 1 class per step. As shown in Tab. 4, compared with the model trained on VOC only, we get 1.34 p.p performance uplift on all metric in PASCAL VOC 15-1 setting, indicating the strong transfer ability of objectness.
Effectiveness of objectness distillation. We ablate objectness distillation components with optimal configuration on other parts. The ablation is conducted in four cases: (1) $\mathcal{L}_{\mathrm{mask-kd}}$ for mask only, (2) $\mathcal{L}_{\mathrm{os-kd}}$ for objectness score only, (3) $\mathcal{L}_{\mathrm{mask-kd}}+\mathcal{L}_{\mathrm{os-kd}}$ for mask and objectness score, and (4) $\mathcal{L}_{\mathrm{mask-kd}}+\mathcal{L}_{\mathrm{os-kd}}+\mathcal{L}_{\mathrm{pe-kd}}$ for mask, objectness score, and position. As shown in Tab. 5, we report the performance in PASCAL VOC 15-1 setting. By comparing case 1 and 3, without $\mathcal{L}_{\mathrm{os-kd}}$, base class mIoU degrades significantly as the base objectness score diminishes. By comparing case 2 and 3, we observe slight forgetting of mask proposals without $\mathcal{L}_{\mathrm{mask-kd}}$ ($\sim$2 p.p drop on all metric), showing the forgetting robustness of mask proposals. The position distillation aims for better alignment of localized information and leads to better classification results, with $\sim$4 p.p all performance gain when comparing case 3 and 4.
Effectiveness of class distillation. We perform ablation in VOC 15-1 setting by analyzing the effectiveness of $\mathcal{L}^{\mathrm{u}}_{\mathrm{cls-kd}}$ and $\mathcal{L}^{\mathrm{m}}_{\mathrm{cls-kd}}$ in Sec. 3.3.2. The results in Tab. 6 indicate $\mathcal{L}^{\mathrm{u}}_{\mathrm{cls-kd}}$ playing a pivotal role in alleviating forgetting, coinciding with pseudo-labeling in previous methods [12, 5]. With $\mathcal{L}^{\mathrm{m}}_{\mathrm{cls-kd}}$, the forgetting issue is mitigated even further.
Table 6: Ablation results on class distillation in VOC 15-1 setting.
$\mathcal{L}^{\mathrm{m}}_{\mathrm{cls-kd}}$ $\mathcal{L}^{\mathrm{u}}_{\mathrm{cls-kd}}$ 1-15 16-20 all avg
34.63 42.36 36.47 53.92
65.24 39.53 59.12 64.23
69.77 43.62 63.54 70.63
Table 7: More ablation results on stage 2 components in PASCAL VOC 15-1 setting, where “TQ” denotes task queries, “Aux” denotes auxiliary loss $\mathcal{L}_{\mathrm{cls-aux}}$, and “Focal” denotes focal loss $\mathcal{L}_{\mathrm{cls}}$.
TQ Aux Focal 1-15 16-20 all avg
59.30 25.96 51.36 59.44
63.66 29.32 55.48 62.45
65.21 35.53 58.14 63.93
66.43 40.20 60.18 64.72
67.91 38.41 60.89 65.06
69.77 43.62 63.54 70.63
Effectiveness of stage 2 other components. As shown in Tab. 7, under optimal distillation configurations, we investigate the contribution of stage 2 other components, including task queries (TQ), focal loss $\mathcal{L}_{\mathrm{cls}}$ (Focal) in Sec. 3.2.2, and auxiliary loss $\mathcal{L}_{\mathrm{cls-aux}}$ (Aux) in Sec. 3.3.2. Note that when not using task queries, one query is used across all tasks; when not using focal loss, cross-entropy is applied. We first ablate the improvements with each improvement only. First, with task queries, we get $\sim$4 p.p improvement on all metric as they reduce the learning interference between tasks. Second, the auxiliary loss assists in learning without forgetting new classes similar to old classes by increasing $\sim$7 p.p of all metric. Third, when using the focal loss to smooth the class probability distribution, it facilitates the distillation process and results in $\sim$9 p.p mIoU gain. With all the improvements, we boost the final performance by $\sim$12 p.p.
5 Conclusion
In this paper, we present CoMasTRe, a continual segmentation framework by disentangling the challenging segmentation problem to objectness learning and class recognition. To leverage the properties of objectness, we propose a two-stage query-based segmenter and distill objectness and classification knowledge separately to alleviate forgetting. Extensive experiments show that our method achieves considerably better performance compared with state-of-the-art methods on PASCAL VOC 2012 and ADE20K. For future works, CoMasTRe provides a decoupled way to tackle continual semantic segmentation and could be extended to continual panoptic and instance segmentation.
Acknowledgments. This work was supported by the National Key R&D Program of China (No. 2022YFE0200300), the National Natural Science Foundation of China (No. 61972323, 62331003), Suzhou Basic Research Program (SYG202316) and XJTLU REF-22-01-010, XJTLU AI University Research Centre, Jiangsu Province Engineering Research Centre of Data Science and Cognitive Computation at XJTLU and SIP AI innovation platform (YZCXPT2022103).
Supplementary Material
This supplementary material first presents a detailed workflow of the class recognition process. Next, it provides more comparisons with state-of-the-art methods. Finally, more ablation studies and detailed results are reported.
Appendix A Workflow of Stage 2 Class Recognition
In detail, we present the class decoder architecture in Fig. A1. It is a single Transformer decoder block whose key (K) and value (V) are pixel embedding; query (Q) is positional embedding with task query. Note that only one positional embedding (red), together with one task query (blue), is fed through the class decoder once a time. The task embedding (purple) is the corresponding output of the task query, and the other embedding (gray) is discarded.
Figure A1: Class decoder architecture.
We also illustrate the stage 2 class recognition process in Algorithm A1. The class decoding process requires positional embeddings, task queries, and pixel embeddings as input and outputs the class probability of all proposals.
Algorithm A1 Class recognition at step $t$
$\mathcal{E}^{\mathrm{m}}_{\mathrm{pos}}=\{e_{1},\dots,e_{M}\}$: matched positional embeds
1:$\mathcal{Q}_{\mathrm{task}}^{t}=\{\theta_{1},\dots,\theta_{t}\}$: task queries at step $t$
2:$\Phi^{t}=\{\phi_{1},\dots,\phi_{t}\}$: classifiers at step $t$
[…]
5:$\mathcal{Q}_{\mathrm{cls}}^{i,j}\leftarrow(e_{i},\theta_{j})$
6:$k_{i,j}\leftarrow g(\mathcal{Q}_{\mathrm{cls}}^{i,j},\mathcal{E}^{l}_{\mathrm{pixel}})$
7:$z_{i,j}\leftarrow\phi_{j}(k_{i,j})$EndFor
8:$z_{i}\leftarrow[z_{i,1},\dots,z_{i,t}]$
9:$p_{i}\leftarrow\mathrm{sigmoid}(z_{i})$EndFor
10:$\mathcal{P}^{t}\leftarrow\{p_{1},\dots,p_{M}\}$
[…]
Appendix B More Comparisons with State-of-the-arts
We compare CoMasTRe with state-of-the-art continual segmentation methods. Here, we additionally report quantitative results in ADE20K 50-50 setting and qualitative results of PASCAL VOC 15-1 and ADE20K 100-10.
Quantitative results in ADE20K 50-50 setting. As shown in Tab. A1, we report benchmark results in ADE20K [58] 50-50 setting. The results show that we achieve a new state-of-the-art in this setting with 0.32 percent point (p.p) leap on all metric compared with CoMFormer [5]. Specifically, our method reaches a competitive new class performance while maintaining better old knowledge (0.63 p.p gain compared with CoMFormer). Furthermore, CoMasTRe performs better across all learning steps, surpassing CoMFormer by 3.12 p.p and even can beat the highest per-pixel method PLOP [12] by 0.30 p.p on avg.
Table A1: Benchmark results in ADE20K 50-50 setting. The 1st and 2nd highest results are marked in bold and underline.
Paradigm Method 50-50 (3 tasks)
1-50 51-150 all avg
Per-Pixel MiB [3] 45.57 21.01 29.31 38.98
SDR [30] 45.66 18.76 27.85 34.25
PLOP [12] 48.83 20.99 30.40 39.42
REMINDER [34] 47.11 20.35 29.39 39.26
RCIL [53] 48.30 25.00 32.50
Joint 51.21 32.77 39.00
Query CoMFormer [5] 49.20 26.60 34.10 36.60
Joint 53.40 38.00 43.10
CoMasTRe (ours) 49.83 26.56 34.42 39.72
Joint 54.09 39.49 44.36
Figure A2: Qualitative results compared with CoMFormer [5] in PASCAL VOC 15-1 setting.
Qualitative results on PASCAL VOC 2012. Compared with CoMFormer [5], we visualize the segmentation results under PASCAL VOC 15-1 in Fig. A2. By comparison, our method is more resistant to forgetting old similar classes. For example, CoMFormer starts to forget the horse after learning sheep class at step 2 (first row) and misrecognizes the dog as a sheep at step 4 (third row). Additionally, our method is less prone to overconfidence on new classes, e.g., CoMFormer falsely recognizes a sofa at step 5 (first row), but our method does not.
Qualitative results on ADE20K. As shown in Fig. A3, we visualize the results in ADE20K 100-10 settings. The visualization suggests that our method preserves better knowledge of previous classes. Compared with CoMFormer [5], our method correctly segments the field in the first row, the transporter in the second row, the mirror in the third row, the mountain in the fifth row, and the house in the sixth row. In addition, CoMasTRe also proposes better masks for old classes. For example, our method maintains the knowledge of the door in the fourth row (yellow box) and the boat in the sixth row (red boxes).
Appendix C More Ablation Studies and Detailed Results
We present more ablation studies on the forgetting of objectness and the effectiveness of objectness score reweighting. In addition, per-class segmentation results on PASCAL VOC are reported.
Table A2: Average recall (AR) of mask proposals in PASCAL VOC 15-1 setting, where $\mathcal{L}_{\mathrm{mask-kd}}$ for mask distillation, $\mathcal{L}_{\mathrm{os-kd}}$ for objectness score distillation, $\mathcal{L}_{\mathrm{pe-kd}}$ for position distillation, $s$ for objectness score, and $\alpha$ for high objectness threshold.
# Case AR
$s>0$ $s>\alpha$
1 $\mathcal{L}_{\mathrm{mask-kd}}$ 55.89 13.75
2 $\mathcal{L}_{\mathrm{os-kd}}$ 50.92 47.84
3 $\mathcal{L}_{\mathrm{mask-kd}}+\mathcal{L}_{\mathrm{os-kd}}$ 55.63 49.02
4 $\mathcal{L}_{\mathrm{mask-kd}}+\mathcal{L}_{\mathrm{os-kd}}+\mathcal{L}_{\mathrm{pe-kd}}$ 55.84 48.83
Objectness forgetting analysis. We analyze the forgetting of objectness by ablating objectness distillation components. The ablation is conducted in the same cases as in Tab. 5 in the main text. For each case, we use average recall (AR) to indicate the performance of mask proposals. Here, $s$ stands for the objectness score, and $\alpha$ is the objectness threshold during inference. As shown in Tab. A2, we report the performance in PASCAL VOC [14] 15-1 setting. By comparing case 1 and 3, without $\mathcal{L}_{\mathrm{os-kd}}$, AR ($s>0$) remains unchanged, but AR ($s>\alpha$) diminishes (-35.27 p.p AR), which means the objectness scores fail to indicate old class objectness. By comparing case 2 and 3, we observe slight forgetting of mask proposals without $\mathcal{L}_{\mathrm{mask-kd}}$ (-4.71 p.p AR), showing the forgetting robustness of mask proposals. When comparing case 3 and 4, we find position distillation contributes most to continual classification (see Tab. 5 in the main text), as the AR changes are negligible (underlined in Tab. A2).
Table A3: Ablation results of objectness score reweighting in PASCAL VOC 15-1 setting by varying reweighting strength $\beta$. Note that no reweight is applied when $\beta=0.0$.
# $\beta$ 1-15 16-20 all avg
1 0.0 67.47 41.83 61.37 68.47
2 1.0 69.79 42.98 63.41 70.35
3 2.0 69.77 43.62 63.54 70.63
Figure A3: Qualitative results compared with CoMFormer [5] in ADE20K 100-10 setting.
Table A4: Per-class segmentation results on PASCAL VOC 2012 in mIoU (%). Incremented classes (inc) are marked in green.
Setting bg aero bike bird boat bottle bus car cat chair cow table dog horse mbike person plant sheep sofa train tv base inc all
19-1 (2 steps) 93.7 91.9 42.8 88.9 65.4 81.1 87.9 80.7 91.9 37.0 80.1 51.3 86.0 82.9 82.0 87.0 56.6 87.3 50.6 77.4 69.5 75.1 69.5 74.9
15-5 (2 steps) 93.5 90.4 43.6 90.8 64.2 82.8 88.2 88.6 94.1 42.8 80.9 70.5 89.4 82.7 84.6 88.4 39.4 59.1 38.9 62.0 60.2 79.7 51.9 73.1
15-1 (6 steps) 88.9 86.4 38.0 82.6 53.2 76.8 76.8 83.2 82.5 36.2 59.3 48.4 80.4 63.1 74.7 85.8 29.9 47.1 33.0 55.4 52.7 69.8 43.6 63.5
Joint 94.3 91.8 43.0 91.1 65.3 85.2 92.2 87.2 93.1 44.2 85.3 69.4 90.5 86.9 85.3 88.4 60.7 83.9 48.1 85.4 68.3 78.1
Effectiveness of objectness score reweighting. In Tab. A3, we ablate the effectiveness of objectness score reweighting mentioned in Sec. 3.3.1 of the main text. The reweight strength $\beta$ is set to 0.0, 1.0, and 2.0, respectively. Please note that when $\beta$ is set to 0.0, the distillation is equivalent to the regular unweighted one. The results show that the reweighting ($\beta=2.0$) leads to 2.17 p.p performance gain on all metric compared with the unweighted version.
Per-class results on PASCAL VOC. In Tab. A4, we provide per-class experimental results on PASCAL VOC 2012 in different continual segmentation settings. The results indicate that in addition to learning new classes, the old class performance can be further strengthened in later learning steps compared with joint baseline, such as sheep class in 19-1 (+3.4 p.p) and car class in 15-5 (+1.4 p.p).